OpenStudy (anonymous):

A cargo plane left an airport at noon and flew toward New York at the average rate of 300 mph. at 2pm a jet plane left the same airport for New York and flew the same route as the cargo plane at the average rate of 500mph. how many miles did the jet plane fly before it overtook the cargo plane?

OpenStudy (anonymous):

\[600+300t=500t\] solve for t to get the time they met

OpenStudy (anonymous):

reasoning is that in two hours the first plane went 600 miles, and then its distance is \(600+300t\) and the other planes distance is \(500t\) since they meet at the same distance can set \(600+300t=500t\) and solve for t to get \(t=3\)

OpenStudy (anonymous):

in 3 hours both planes have gone 1500 miles

OpenStudy (hoblos):

let t= time that the jet plane fly before it overtook the cargo plane 500t = 300(t+2) 500t = 300t +600 200t= 600 t=3 hr distance = speed x time = 500 x 3 =1500 mile

OpenStudy (anonymous):

thanks..!!